Staybrite Plc has just released its financial results for 2018. It achieved sales of £300m, and forecast that these were expected to grow at 6% per year.
Staybrite Plc currently has cash on deposit of £45m, debt of £15m and has issued 18m shares. It is taxed at a rate of 28%.

You estimate the following information:
EBIT is 6% of sales
Increases in both net investment and in net working capital will be 7% of any increase in sales. The firm’s WACC is expected to be 10%.
Using a free cash flow approach, calculate a fair price for staybrite Plc shares.
Answer #1

Given:

Growth in sales = 6% per year

EBIT = 6% of Sales

WACC = 10%

Increase in Net Investment and Net Working Capital = 7% of increase in sales

Number of shares = 18 million

Debt = £ 15 m

Calculation of Free Cash Flow to the firm :

S.No. Particulars a. Sales b. EBIT (6% *a) C. Tax A 28% (b*28%) d Increase in Net Investment & Net Working Capital e. Free Ca

Value of The firm = Free Cash Flow for Projected 2019/Wacc-Growth Rate of Free Cash Flow

= 12.4776/0.10-0.06

= 12.4776/0.04

= £ 311.94 m

Value for Equity Share Holders = Value of the firm - Value of Debt

= £ 311.94 m - £ 15m

= 296.94

Fair Price for Staybrite Plc Shares = 296.94/18 m shares

= £ 16.50

Note:

1. In the absence of information about depreciation & amortization nothing has been added added back on account of depreciation & amortization though it is a non cash expenditure.

2. Combined increase of 7% on account of Net Investment and Net Working Capital has been considered.